OpenStudy (anonymous):

Hey guys I need help with this problem. Is 1/4 the root of the equation 4/y + 6/2 y =28?

OpenStudy (anonymous):

is that an equation or an expression you are trying to evaluate ?

OpenStudy (anonymous):

Equation

ganeshie8 (ganeshie8):

put y = 1/4 in the equation. if it satisfies the equation, 1/4 is a root. if it doesnt satisfy the equation, its not a root

OpenStudy (anonymous):

My answer is different from the book

OpenStudy (anonymous):

is the equation 4/y + 6/2 = 0 ? if so, then follow ganeshie8's advice and let y = 1/4 to see what you obtain

OpenStudy (anonymous):

Its 6/2 y not 6/2

OpenStudy (anonymous):

4/y + 6/y where y = 28 ? What do you obtain with this substitution ?

ganeshie8 (ganeshie8):

\(\large \frac{4}{y} + \frac{6}{2y} = 28\)

ganeshie8 (ganeshie8):

i think he means that.. ?

OpenStudy (anonymous):

4/y + 6/2 y or 4/y + 6/(2y) ?. In either case, ganeshie8's advice is applicable

OpenStudy (anonymous):

No mimic I know what to do but my answer is different. I got 7/4 as the answer

OpenStudy (anonymous):

Yes ganeshie8 that is the equation

OpenStudy (anonymous):

I see. What steps did you use? Did you set up the equation this way: \[\frac{ 4 }{ y } + \frac{ 6 }{ 2y } = 28\] If so, what did you do next ?

OpenStudy (anonymous):

I did 4* 1/4 + 6/2 * 1/4. I reduced everything and got 1/1 + 3/4 then found the LCD so its 4/4 + 3/4 =7/4

OpenStudy (anonymous):

I see. When you substitute y = 1/4, be careful. Since y is in the denominator, your first step would be: \[\frac{ 4 }{ \frac{ 1 }{ 4 } } = 16\] Can you figure out the next step?

OpenStudy (anonymous):

Wait a minute please

OpenStudy (anonymous):

Yes I got 28 but why didn'tmy first solution work?

OpenStudy (anonymous):

You multiplied by 1/4 instead of dividing by 1/4.

OpenStudy (anonymous):

Anything elseyouwantto add? ?

OpenStudy (anonymous):

BC I'm thinking that the y is applying to everything but iguess that in this case where the why y is important well thank you sir or madame